NEXT is contemplating a T-shirt advertising promotion. Monthly sales data from T-shirt shops
marketing the “Eye Watch NEXT” design indicate that
Q = 1,500 – 200P
Where: Q is T-shirt sales and P is price.
A. How many T-shirts could NEXT sell at $4.50 each?
B. What price would NEXT have to charge to sell 900 T-shirts?
C. At what price would T-shirt sales equal zero?
D. Calculate the point price elasticity of demand at a price of $5?
1
Expert's answer
2017-10-12T02:59:07-0400
Quantity of T-shirts at the price $4.50 is:
Q = 1,500 – 200*4.5 = 600 (T-shirts)

Price for selling 900 T-shirts is:
900 = 1,500 – 200P
200P = 600
P = $3
T-shirts sales equal zero at the price:
1,500 – 200P = 0
200P = 1,500
P = $7.5
Quantity at the price of $5:
Q = 1,500 – 200*5 = 500 (T-shirts)
Price elasticity of demand:
E = (Q2-Q1)/Q1 / (P2-P1)/P1 = (500-600)/600/($5-$4.5)/$4.5 = -1.5